# AI-Powered Episodic Future Thinking

### v1

@@ 3.5.5 Interface
“I didn’t like how hard it was to copy what was said. I had a hard time copying the rewrites.” [P137]
Voice control was also suggested as a potential accessibility feature to simplify interactions:
“Voice control. Make it easier.” [P137]
Beside these, participants found it easy to use [P119], [P120], [P136]. One participant found the interface intuitive and appreciated the straightforward instructions:
“It was easy to input information, and the instructions were clear.” [P124]
@@
4 Discussion
The goal of this study is to explore users’ experiences and perceptions of EFTeacher, an AI chatbot designed for EFT generation. Specifically, we aim to understand users’ overall experiences with the chatbot, their impressions of its functionality, and their perceptions of working with an interactive AI tool for EFT generation.
@@ 4 Discussion
Our analysis indicates that the chatbot’s descriptive language was a polarizing feature among participants. Some individuals appreciated its vivid descriptions, which enhanced their sense of vividness and made their interactions feel richer. Participants highlighted the chatbot’s ability to create detailed visuals and make scenarios feel more tangible. However, others felt that the language was overly elaborate and occasionally unnecessary. For these users, the level of detail made conversations less natural.
@@
Interestingly, this issue could potentially be mitigated through the chatbot’s adaptability features, which allow users to request simpler or less descriptive responses. However, it became evident that participants did not utilize it frequently, suggesting a gap in feature accessibility or motivation to adjust responses mid-interaction. The chatbot’s adaptability was highlighted as a key strength by some partiicpants, allowing them to refine responses and tailor interactions to their preferences. However, some participants did not utilize these features effectively, leading to missed opportunities for improving their experiences. Multiple engagements with the chatbot could facilitate a learning process that helps users become more familiar with these adaptability features, ultimately improving their ability to personalize interactions.
Participants frequently highlighted the positive emotional impact of their interactions with the chatbot, emphasizing its ability to inspire hope, optimism, and validation. The chatbot’s generated EFTs encouraged users to reflect on their goals and aspirations, often guiding them to explore aspects of their future they might not have otherwise considered.
@@
The vividness and personalization of the chatbot’s EFTs played a key role in creating these positive emotional experiences. EFTeacher demonstrated potential in helping users articulate their thoughts with greater detail and structure, encouraging them to engage more deeply with their personal visions. This approach not only strengthened their emotional connection to the conversation but also nurtured a sense of optimism about the future.
This aligns with the core principles of EFT interventions, where individuals vividly imagine and describe positive future events. This process suggests that an AI chatbot could be a valuable tool for effective EFT interventions by encouraging users to consider the future and helping them focus on their long-term goals.
@@
While the reflective nature of the interactions was highly valued, some participants noted that the level of detail could occasionally feel overwhelming and wordy. This stems from AI-generated language, which, coming from a machine, does not always sound natural. However, despite these constraints, AI can still be a valuable tool for structuring and refining EFT-related written content, making it a supportive tool, especially for those who have difficulty with writing. Participants mentioned that the repetition in the chatbot’s responses negatively impacted their interactions. This repetition stemmed from the chatbot prompting users for seven different events at seven different times, following the same process each time. Participants described this as overly repetitive and excessive. Reducing repetitive patterns could improve the chatbot’s ability to deliver more natural, human-like interactions. To improve user experience, our research group has decided to conduct the EFT teacher interview across four (instead of seven) different time frames in future iterations.
5 Design Insights
5.1 Integrating Voice Capabilities into EFTeacher
Based on the findings from our thematic analysis, integrating voice capabilities into EFTeacher could enhance its effectiveness in creating a more engaging and natural conversational experience. Participants consistently praised EFTeacher’s ability to emulate human-like conversation, with its tone being described as thoughtful and lifelike. Introducing voice as a medium of interaction could address these limitations by enhancing the emotional and contextual depth of communication.
@@ 5.1 Integrating Voice Capabilities into EFTeacher
Prior research has supported the importance of voice in human-agent interaction. According to [65], voice impacts not only the content of the speech but also the overall perception of the agent’s personality and context-specific relevance. Additionally, audio delivery has been explored in the context of Episodic Future Thinking (EFT) interventions with promising results. Studies such as [25, 17, 66] support the use of audio in EFT interventions. Participants can generate cues and then vividly imagine or listen to audio recordings of their EFT cues during tasks that assess delay discounting. This approach has demonstrated effectiveness in enhancing the intervention experience by leveraging audio’s ability to evoke richer emotional and cognitive engagement.
A scoping review of the literature further suggests that voice assistants can serve as a valuable tool in the research of behavioral health interventions [67]. This supports the potential for EFTeacher to leverage voice capabilities not only to enhance user engagement but also to support interventions aimed at improving behavioral and emotional well-being.
@@
Voice integration also addresses accessibility concerns raised during usability testing. For example, participants suggested features like voice control to simplify interactions, especially for users who might find typing or reading responses challenging. Voice output and input would not only improve usability but also expand EFTeacher’s accessibility to a broader audience.
5.2 Leveraging Visual Capabilities for EFTeacher
@@
Advancements in LLMs present a promising avenue for building effective human-AI interaction. The use of multimodal features of LLMs is increasingly gaining attention in recent technologies, as demonstrated in [68, 69, 70]. One such feature is the ability of models to generate images from text prompts, as explored in [69]. Integrating this capability into EFTeacher could allow the chatbot to produce visual EFTs in response to user queries, making interactions more engaging and particularly useful for tasks that require visual representations, such as illustrating future events based on user prompts and generating written EFTs.
The benefits of incorporating visual elements are further supported by research on EFT. A study [71] investigated the effects of drawn and written future cues, and found that both drawn and written future cues led to reduced delay discounting compared to the control group, with drawn cues offering unique advantages in terms of personal relevance and the vivid encoding of details. These findings suggest that enabling EFTeacher to generate or support visual representations of future scenarios could further enhance its impact on future-oriented decision-making.
@@
Moreover, incorporating visual elements could address a key finding from our study: while some participants appreciated the AI-generated language, others found it overly elaborate or unnecessary. This discrepancy highlights the need for alternative modes of interaction, such as visualizations, for users who may not connect with the AI chatbot’s language or who prefer a less verbose experience. By offering visual representations alongside text-based descriptions, EFTeacher could accommodate a broader range of user preferences.
6 Limitations and Future work
For this study, we asked participants to describe seven future events. This approach helped us gather more AI-generated EFT data, but made the experience feel repetitive for some participants. In future studies, we plan to reduce this to four events, which we believe will make the interaction more positive and less repetitive while still collecting valuable data.

### v2

@@ 3.5 Assessment Regarding User Perceptions
137 “I didn’t like how hard it was to copy what was said. I had a hard time copying the rewrites.”
139 “larger font sizes would be nice. ”
117 “not giving a word count. If there is an expectation of level of detail or goal then nice to be able to track that from the start"
3.5.1 Human-like dialogue and communication proficiency
Participants reported on EFTeacher’s dialogue, particularly how conversational, natural, and human-like the interactions felt. A central aspect of this perception was around the degree to which interactions felt conversational and human-like. Several noted that the system describing its conversational style as authentic and engaging [P105, P109, P112, P118].
They highlighted EFTeacher’s ability to maintain dialogue that resembled human interaction, using descriptors such as natural [P139], genuine [P137], and referencing features like leading questions [P117]. The chatbot’s responses were also praised for being empathetic and contextually relevant: “It really interacted with me like a human! I expressed so much joy at something it wrote and said I might need to save it to read later, and it responded so thoughtfully.” P[118]
Despite its strengths, some participants found the interactions too polished or scripted, detracting from the perceived naturalness P[104, 120, 125], as participant P[115] said: “I believe the EFTeacher gave you more than an average human in regards to the descriptive words and paragraphs.”
Repetition also emerged as a common critique. While many users appreciated the flow of the conversation, some noted that EFTeacher tended to repeat similar requests for details, making the interaction feel less dynamic P[119]: “Not only was their storytelling great, but it seemed to flow somewhat like a real conversation, except for the repetitiveness of asking for the details it needed each time.” P[123]
These insights suggest that while EFTeacher demonstrates an ability to create engaging and human-like conversations, there is room for improvement. Participants’ feedback highlights a dual expectation: while many users appreciated EFTeacher’s engaging and human-like dialogue, the repetition of similar requests for details caused frustration for some. Striking a balance between maintaining high-quality responses and reducing repetitive patterns is crucial for the refinement of EFTeacher.
This theme aligns with prior work suggesting that perceived naturalness in chatbot interactions enhances user engagement and the efficacy of conversational agents [li2024finding]. Communication that feels more human-like contributes positively to user participation and receptiveness to the intervention [olafsson2020motivating]. Tailoring a chatbot’s persona and communication style to user preferences, and to create a stronger sense of connection, increases the likelihood of sustained engagement [li2023exploring].
3.5.2 Adaptability in response to participants’ input
Participants described how EFTeacher adapted its responses based on their preferences, enabling them to make edits, request word changes, and engage in flexible, personalized interactions. Participants appreciated EFTeacher’s responsiveness to real-time feedback, especially when editing specific words or rephrasing content, and found it easy and intuitive to request changes or content rewrites P[116, 118, 139]: “It was very easy to have EFTeacher make changes. I didn’t know if I could ask for specific words to be changed, so I just tried it, and it worked! When I asked for less vivid wording, it changed it to something that I liked better.” P[112]
Participants noted that EFTeacher not only accepted edits but also clarified intent by paraphrasing their input or identifying when an input was incomplete P[109]: “I liked the way it confirmed what you were trying to convey. It was great that EFTeacher understood my words and paraphrased it back to me.” P[115].
They highlighted that EFTeacher felt more grounded and tailored to the user compared to other AI systems P[105, 120].
However, one participant identified limitations in EFTeacher’s ability to fully simplify responses when requested. While the tool was adaptable, certain adjustments fell short of expectations: “Even when asked to simplify it was still more than I thought a simple description would be. For some people less words work better for their understanding. So when asked to simplify something, I would expect a shorter answer with right-to-the-point statements.” P[125]
Most of the feedback highlights the significance of personalization and adaptability in improving user interaction, emphasizing the importance of addressing diverse user expectations. Prior study [li2024finding] supports this by showing that users highly value the ability to tailor their AI chatbots to their preferences. This customization includes both the chatbot’s visual aspects and its conversational style, reflecting a strong user desire for personalized AI experiences. The present findings underscore the need for adaptive systems when responding to user intent, for user-centered AI systems.
3.5.3 Imagination facilitated through future-oriented prompts
EFTeacher supported users in visualizing future events by prompting imaginative thinking. Participants described EFTeacher as a powerful tool for guiding imagination and enhancing their ability to envision detailed, future-oriented scenarios P[127, 123]. They described it as enhancing their milestones and making them feel even more meaningful P[120], painting a vivid picture P[125], and triggering their imagination to visualize these moments P[122]. Many noted helped them articulate life goals and future plans in ways they hadn’t considered before.
The language was often described as descriptive, and the use of descriptive wording transformed abstract thoughts into imagery stories P[122, 115, 104], that participants would not have generated on their own. As one participant shared: “Used lots of adjectives. The statement by the AI describes things better than I could put it. It helped me imagine actually being there.” P[125]
Several participants highlighted that the guided prompts and questions posed by EFTeacher were particularly effective in facilitating the imagination of their stories. The questions were described as helpful P[109, 112], well-designed P[103], thought-provoking P[119], and detail-oriented P[137] for storytelling: “I liked how it asked questions about my storyline because it made me think about the setting, people, and everything. It really helped bring out the writer in me.” P[122]
EFTeacher also helped participants who struggled to formulate ideas, particularly for distant future scenarios. The tool’s prompts provided structure and inspiration, helping users transform ideas into narratives P[127, 123, 106, 137], and through the added details P[106,138]: “I was really struggling to imagine events ten years from now, and it helped me string together some of my piecemeal thoughts!” P[118]
However, one participant mentioned expecting the chatbot to take the lead in storytelling or in providing meaningful new ideas beyond their initial input: “It felt like I was the one coming up with every story, and they were just riffing on what I came up with by myself. I don’t think this is bad, but I didn’t feel like it was giving me anything to go off of besides the timeframe the story took place in.” P[123]
The ability of EFTeacher to stimulate the imagination through vivid descriptions, questions, and prompts is a standout feature. It has the potential to facilitate imagination, making it an effective tool for users to envision and articulate their future.
3.5.4 Reflection on using EFTeacher and takeaways from the interaction with EFTeacher
Participants shared how EFTeacher prompted them to engage in meaningful reflection, fostering emotional insight, future-oriented thinking, and personal awareness. Participants reported that interacting with the system helped them articulate and visualize goals, explore new aspirations, and experience emotional feelings.
Many participants described how EFTeacher encouraged them to reflect on their life direction, consider long-term goals, and think beyond the present moment. The tool provided structured prompts that initiated future-oriented thought P[122, 106, 137, 115]: “Don’t typically think very far into the future, and it made me consider where I want my life to be in X months/years.” P[106]
Participants reported that the emotionally rich and vivid content generated by EFTeacher created feelings of excitement P[138] and hope P[107]. Seeing their imagined futures articulated helped some feel encouraged about what lies ahead P[120].
The added details and language generated with EFTeacher helped the participants construct emotional narratives P[115, 125] about their future. They noted that the system helped them articulate their thoughts in ways that improved clarity and storytelling: “It was nice that it suggested more helpful details in my writing. It kept me focused and thinking about my future.” P[137]
EFTeacher facilitated personal reflection by encouraging participants to explore future goals and possibilities they had not previously considered. Its vivid and emotionally engaging prompts led users to feel hope, excitement, and validation, helping them reflect on life direction and envision meaningful future scenarios. This aligns with prior research showing how AI tools can support introspection. Similar to AI-driven journaling platforms [nepal2024contextual], where carefully designed prompts are used to elicit self-reflection and emotional processing, EFTeacher’s personalized and vivid prompts guided users to articulate their aspirations and feelings. Similarly, Liu et al. [kocielnik2018designing] demonstrate that conversational agents can support reflection in daily life through structured prompts. EFTeacher thus serves not only as an EFT cue generator but also as a tool for self-reflection and future-oriented thinking.
3.5.5 Feedback about interface, visual design and ease of interaction
Participants shared a range of impressions about EFTeacher’s interface, regarding its usability and visual design. Overall, many found the interface intuitive and easy to use. For instance, one participant noted, “It was easy to input information, and the instructions were clear.” P[124]. Others echoed this sentiment, describing the platform as simple to navigate and accessible, especially when submitting responses P[119, 120, 136].
At the same time, participants identified specific areas for improvement that impacted readability, accessibility, and overall user experience. A common concern was the need for better visual clarity, particularly around color contrast P[129] and font P[105]. Others proposed enhancements to make content more accessible, such as audio support P[137] or voice control: “No option for audio and visual. Hearing the descriptions and scenarios aloud would have added good detail.” P[129].
Usability suggestions focused on improving interaction features that could support task completion. For example, users requested a word count to help guide their responses P[117]. Another participant noted difficulty in copying and reusing texts P[137].
@@ 3.5.5 Feedback about interface, visual design and ease of interaction
3.6 EFT Cue Text Analysis
We analyzed several linguistic features of the EFT cues generated with EFTeacher for 25 participants, across seven different timeframes, resulting in a total of 175 cue texts. Table 3 summarizes descriptive statistics for some textual features extracted from cue texts, including: number of words and unique words, and the Type-Token Ratio (TTR), which quantifies lexical diversity by dividing the number of unique words by the total word count. These metrics were computed individually for each cue and then averaged [lexicalrichness]. Flesch reading score and Flesch-Kincaid grade level were also computed [textstat]. The Flesch reading score reflects how easy a text is to read, while the Flesch-Kincaid grade estimates the U.S. school grade level required to understand the text. Sentiment polarity and subjectivity were also determined [textblob]. Polarity ranges from -1 (negative) to 1 (positive), and subjectivity from 0 (objective) to 1 (subjective).
Table 3: Statistics for Cue Text Features
Feature Mean Median Std Dev Min Max IQR
Number of words 108.98 106.00 25.31 41.00 220.00 31.00
Number of unique words 76.27 75.00 15.33 35.00 135.00 18.50
TTR 0.71 0.71 0.04 0.57 0.85 0.06
Flesch readability score 58.74 59.03 9.44 35.01 82.48 12.75
Flesch-Kincaid grade level 10.04 9.87 1.92 5.04 15.35 2.60
Sentiment polarity 0.33 0.32 0.14 -0.08 0.82 0.19
Sentiment subjectivity 0.57 0.58 0.11 0.22 0.83 0.13
According to the table, the average number of words per cue was 108.98, with a median of 106. Cue lengths ranged from 41 to 220 words, indicating a moderately wide variation in text length. The interquartile range (IQR) was 31, suggesting that most cues fall within a relatively consistent length range. The average number of unique words per cue was 76.27, with a median of 75 and values ranging from 35 to 135. The standard deviation of 15.33 indicates a moderate degree of variation in vocabulary diversity across cues—some participants used relatively limited vocabularies, while others used a broader range of terms.
The type-token ratio (TTR), a commonly used measure of lexical diversity, was calculated as the ratio of unique words to total words for each cue and then averaged across texts. The mean TTR was 0.71, with a standard deviation of 0.04 and a narrow interquartile range (IQR) of 0.06, suggesting that lexical diversity was relatively stable across cues. A higher TTR indicates a larger proportion of unique words, reflecting more varied language use.
The average Flesch Reading Ease score across cue texts is 58.74, with a median of 59.03, a standard deviation of 9.44, an interquartile range (IQR) of 12.75, and a range from 35.01 to 82.48. The standard deviation and IQR indicate a moderate spread, suggesting that while most cues fall near the average readability level, some are significantly easier or harder to read. A score around 59 falls into the “fairly difficult to read” category [textstat]. The average Flesch-Kincaid Grade Level is 10.04, with a median of 9.87, a standard deviation of 1.92, and an IQR of 2.60, suggesting that most texts are suitable for readers at the 10th to 12th grade level, with relatively low variability across cues [textstat]. The average sentiment polarity across cue texts is 0.33, with a median of 0.32, a standard deviation of 0.14, an interquartile range (IQR) of 0.19, and a range from –0.08 to 0.82. The positive mean and median polarity suggest that, on average, cue texts express a positive emotional tone. Importantly, the minimum polarity of –0.08 is very close to zero, indicating that negative sentiment is not observed; even the least positive cues are nearly neutral in tone. Overall the sentiment remains on the positive or neutral side. The average subjectivity score is 0.57, with a median of 0.58, a standard deviation of 0.11, and an IQR of 0.13, ranging from 0.22 to 0.83. This indicates that the cues are generally moderately subjective, meaning they include personal opinions or affective language rather than purely factual content.
4 Discussion and Design Recommendation
@@
The goal of this study is to explore users’ experiences and perceptions of EFTeacher, an AI chatbot designed for EFT generation. Specifically, we aim to understand users’ overall experiences with the chatbot, their impressions of its functionality, and their perceptions of working with an interactive AI tool for EFT generation. Prior research has shown that the vividness of EFT cues—measured through self-report [brown2022putting] and content characteristics measurement on a Likert scale [rung2020translating]—is associated with reductions in delay discounting. We capture vividness through participant ratings of each cue’s vividness, liking, excitement, and importance (Section 3.4), and through an open-ended survey (Section 3.5). This allowed us to explore not only whether users experienced cues as vivid, but also when vividness was perceived as helpful (e.g., supporting imagination and motivation) or overwhelming (e.g., excessive verbosity). Our thematic analysis indicates that the chatbot’s descriptive language was a polarizing feature among participants. Most individuals appreciated its vivid descriptions, which enhanced their sense of vividness and made their interactions feel richer. Participants highlighted the chatbot’s ability to create detailed visuals and make scenarios feel more tangible. However, others felt that the language was overly elaborate and occasionally unnecessary. For these users, the level of detail made conversations less natural.
@@
Interestingly, this issue could potentially be mitigated through the chatbot’s adaptability features, which allow users to request simpler or less descriptive responses. Indeed, the chatbot’s adaptability was highlighted as a key strength by some participants, allowing them to refine responses and tailor interactions to their preferences. However, some participants did not utilize these features effectively, leading to missed opportunities for improving their experiences. This demonstrated a gap in feature accessibility or motivation to adjust responses mid-interaction. Multiple engagements with the chatbot could facilitate a learning process that helps users become more familiar with these adaptability features, ultimately improving their ability to personalize interactions.
Another notable finding from our thematic analysis is EFTeacher’s ability to facilitate imagination through future-oriented prompts. Participants described how the chatbot’s vivid, descriptive language and structured guidance helped them envision detailed future scenarios, even when initially struggling to generate ideas. Our findings suggest that AI chatbots like EFTeacher can serve not only as tools for cue creation but also as catalysts for future-oriented imagination, helping users simulate possible futures in a structured yet emotionally engaging way.
@@
Participants highlighted the positive emotional impact of their interactions with the chatbot, emphasizing its ability to inspire hope, optimism, and validation. The chatbot’s generated EFTs encouraged users to reflect on their goals and aspirations, often guiding them to explore aspects of their future they might not have otherwise considered. The vividness and personalization of the chatbot’s EFTs played a key role in creating these positive emotional experiences. EFTeacher demonstrated potential in helping users articulate their thoughts with greater detail and structure, encouraging them to engage more deeply with their personal visions. This approach not only strengthened their emotional connection to the conversation but also nurtured a sense of optimism about the future. This aligns with the core principles of EFT interventions, where individuals vividly imagine and describe positive future events. This process suggests that an AI chatbot could be a valuable tool for effective EFT interventions by encouraging users to consider the future and helping them focus on their long-term goals.
@@
While the reflective nature of the interactions was highly valued, some participants noted that the level of detail could occasionally feel overwhelming and wordy. This stems from AI-generated language, which, coming from a machine, does not always sound natural. However, despite these constraints, AI can still be a valuable tool for structuring and refining EFT-related written content, making it a supportive tool, especially for those who have difficulty with writing.
Participants mentioned that the repetition in the chatbot’s responses negatively impacted their interactions. This repetition stemmed from the chatbot prompting users for seven different events at seven different times, following the same process each time. Reducing repetitive patterns could improve the chatbot’s ability to deliver more natural, human-like interactions. To improve user experience, our research group therefore decided to conduct the EFT teacher interview across four (instead of seven) different time frames in future iterations.
Through linguistic feature analysis of cue texts generated by EFTeacher, we found that lexical diversity was high, with a mean Type-Token Ratio (TTR) of 0.71 and an average of 76 unique words per cue. These values indicate a broad range of vocabulary used for the descriptive and detailed language observed in participant feedback. The mean Flesch Reading Ease score was 58.74, classified as “fairly difficult to read” and the average Flesch-Kincaid Grade Level was 10.04, suggesting that the language used may be advanced or verbose for some users. Sentiment analysis revealed a mean polarity of 0.33 (positive) and a subjectivity score of 0.57, indicating that the cues were generally optimistic and included subjective or narrative-style content rather than objective facts.
4.1 Design Recommendations
Participants in the study made suggestions that might lead to a next generation version of EFTeacher. While more development would be needed for such, it is valuable that our study provided a set of design requirements to consider.
@@ 4.1 Design Recommendations
Integrating Voice Capabilities into EFTeacher: Based on the findings from our thematic analysis, integrating voice capabilities into EFTeacher could enhance its effectiveness in creating a more engaging and natural conversational experience. Participants praised EFTeacher’s ability to emulate human-like conversation, with its tone being described as thoughtful and lifelike. Introducing voice as a medium of interaction could address any limitations by enhancing the emotional and contextual depth of communication. Prior research has supported the importance of voice in human-agent interaction. According to [seaborn2021voice], voice impacts not only the content of the speech but also the overall perception of the agent’s personality and context-specific relevance. Additionally, audio delivery has been explored in the context of Episodic Future Thinking (EFT) interventions with promising results. Studies such as [sze2015web, daniel2013future, stein2016unstuck] support the use of audio in EFT interventions. Participants can generate cues and then vividly imagine or listen to audio recordings of their EFT cues during tasks that assess delay discounting. This approach has demonstrated effectiveness in enhancing the intervention experience by leveraging audio’s ability to evoke richer emotional and cognitive engagement.
@@
A scoping review of the literature further suggests that voice assistants can serve as a valuable tool in the research on behavioral health interventions [sezgin2020scoping]. This supports the potential for EFTeacher to leverage voice capabilities not only to enhance user engagement but also to support interventions aimed at improving behavioral and emotional well-being. Voice integration also addresses accessibility concerns raised during usability testing. For example, participants suggested features like voice control to simplify interactions, especially for users who might find typing or reading responses challenging. Voice output and input would not only improve usability but also expand EFTeacher’s accessibility to a broader audience.
@@
Leveraging Visual Capabilities for EFTeacher: Advancements in LLMs present a promising avenue for building effective human-AI interaction. The use of multimodal features of LLMs is increasingly gaining attention in recent technologies, as demonstrated in [pereira2024human, lin2024jigsaw, alalyani2024multimodal]. One such feature is the ability of models to generate images from text prompts, as explored in [lin2024jigsaw]. Integrating this capability into EFTeacher could allow the chatbot to produce visual EFTs in response to user queries, making interactions more engaging and particularly useful for tasks that require visual representations, such as illustrating future events.
@@
The benefits of incorporating visual elements are further supported by research on EFT. A study [carr2021written] investigated the effects of drawn and written future cues, and found that both drawn and written future cues led to reduced delay discounting compared to the control group, with drawn cues offering unique advantages in terms of personal relevance and the vivid encoding of details. These findings suggest that enabling EFTeacher to generate or support visual representations of future scenarios could further enhance its impact on future-oriented decision-making. Moreover, incorporating visual elements could address a key finding from our study: while some participants appreciated the AI-generated language, others found it overly elaborate or unnecessary. This discrepancy highlights the need for alternative modes of interaction, such as visualizations, for users who may not connect with the AI chatbot’s language or who prefer a less verbose experience. By offering visual representations alongside text-based descriptions, EFTeacher could accommodate a broader range of user preferences.
5 Limitations, Conclusions, and Follow-on Work